0.205 grams of molecular compound is dissolved in 110ml of pure distilled water at 25c. The solution freezes point temp is then experimentally measured to be -0.0956C. What is the molecular weight of a compound?


1
Expert's answer
2021-02-09T04:43:17-0500

"\Delta Tf" = Kf×m

"\Delta Tf" = Kf × (w/m)×(1000/V)

-0.0956 = 1.86×(0.205/m)×(1000/110)

m = 32.25 g